how do we do cofunctions?

If sec theta = 3 find the exact value of :
a) cos theta
b) tan^2 theta
c) csc(90 degree - theta)
d) sin^2 theta

Answer by lwsshak3(11628)

If sec theta = 3 find the exact value of :
a) cos theta
b) tan^2 theta
c) csc(90 degree - theta)
d) sin^2 theta
**
for [0,2π]
sec theta=hypotenuse/adjacent side=3/1
opposite side=√(3^2-1^2)=√(9-1)=√8
a) cos theta=1/sec theta=1/3
b) tan^2 theta
tan theta=opp side/adj side=√8/1=√8
tan^2 theta=8
c) csc(90 degree - theta)=sec theta
d) sin^2 theta
sin theta=opp side/hypotenuse=√8/3
sin^2 theta=8/9